# V Holešovičkách 2, 18000 Praha 8, Czech Republic

### v1

@@ 5.1 Relevance of high-eccentricity results
The high-eccentricity and high-convergence-parameter region of parameter space mapped in Fig. 10 and described in Sect. 4.4.3 is the domain not only of the seven-image lens configurations, but also of bulge-shaped critical curves and other peculiarities not occurring at lower eccentricities. These properties of the ellipsoidal NFW lens could not be discovered from low-eccentricity approximations of the lens model, such as those derived from an elliptically distorted lens potential (e.g., Gomer et al., 2023). On the other hand, the astrophysical relevance of such parameter values should be considered.
@@ 5.2 Role of source redshift
The convergence parameter $\kappa_{\text{s}}$ is directly proportional to the angular-diameter-distance ratio $D_{\text{ls}}/D_{\text{s}}$, which increases with source redshift at a rate that depends on the cosmological model (see, e.g., Umetsu, 2020). While some of the presented results such as those in Fig.s 26 or Fig. 13 are independent of $\kappa_{\text{s}}$ and thus of the source redshift, other are presented for specific values of $\kappa_{\text{s}}$, such as those in Fig. 7, Fig. 9, Fig. 15, or Figs. 1718. In the parameter-space maps in Fig. 8 and Fig. 10, different source-plane distances behind a given lens correspond to points along horizontal line segments. These may intersect the plotted boundaries, leading to significant changes in the critical curves and caustics as a function of source redshift.
The dependence of the convergence–shear diagrams on the source redshift is illustrated by the dependence on $\kappa_{\text{s}}$ in Fig. 16. The Jacobian-factor contour plots described in Appendix C (with examples in Fig.s 1920) present an excellent visualization of the variation of the critical curves with source redshift. Since different values of $\kappa_{\text{s}}$ correspond to different nested contours in the same plot, the critical curves for a given range of source redshifts can be directly seen as the corresponding sequence of neighboring contours. For a given lens at a redshift $z_{\text{l}}$, the contours could just as well be labeled by the source redshift $z_{\text{s}}$, with $z_{\text{s}}=z_{\text{l}}$ corresponding to the point at the origin of the plots and higher $z_{\text{s}}$ corresponding to larger contours.
@@ 5.3 Comparison with non-singular isothermal ellipsoid
The comparison in the previous paragraph is relevant to the low-convergence-parameter regime of the ellipsoidal NFW model discussed in Sect. 4.4.2 and mapped in the left panel of Fig. 8. The additional boundaries and lensing regimes appearing at higher $\kappa_{\text{s}}$, discussed in Sect. 4.4.3 and mapped in the right panel of Fig. 8 and in Fig. 10, do not occur in the NIE model. A more detailed comparison of these or other ellipsoidal models is beyond the scope of the present work.
5.4 Possible modifications of model
@@
While the derived analytic formulae for the light deflection by the ellipsoidal NFW model present a large step toward a more astrophysically relevant model from the spherical or ellipsoidal-potential NFW models, structure-formation simulations indicate the potential further steps toward more realistic models. One would be the possible introduction of a core to eliminate the central density divergence. Another would be to relax the homoeoidality of the model: simulations of dark matter halos such as those by Jing & Suto (2002) indicate that their nested constant-density surfaces do not preserve the same ellipsoidal shape. Rather, the axis ratios (here parameterized by the eccentricities $e_{1}$ and $e_{2}$) and axis orientations tend to vary from the center outward. However, such refinements would most probably stay in the realm of numerical modelling and their verification would more likely than not require uniquely precise observational data.
@@ 6 Summary
We applied the described approach in Sect. 4 to study gravitational lensing by a triaxially ellipsoidal dark-matter halo with a NFW density profile (Navarro et al., 1996). The form of the convergence profile given by Eq. (21) as a function of semi-major axis is analogous to the convergence of a spherical NFW halo. However, the convergence parameter depends on the halo shape and orientation, as seen in Eq. (22). The main results of the present work are the analytic formulae derived in Sect. 4.2. The deflection angle for the studied ellipsoidal NFW halo is given by Eq. (25) in terms of the complex scattering function $I$. The components of the reduced deflection angle $\bm{\alpha}_{0}$ are presented in Eq.s (30)–(31) in terms of real functions. These exact closed-form expressions are valid for arbitrary eccentricity $e\in[0,1]$ of the surface-density profile.
@@
The variation of the linear dimensions and enclosed areas of the critical curves and caustics with eccentricity $e$ is demonstrated in Fig. 12 and Fig. 14, respectively. A comparison of the top and bottom rows shows there is a substantial change from a low-convergence to a higher-convergence regime. For low values of $\kappa_{\text{s}}$: with increasing $e$ the tangential critical curve grows prominently horizontally and its area initially declines before increasing to a high-eccentricity peak; the radial critical curve is elongated vertically and its vertical extent initially grows with $e$ before reaching the umbilic transition; the tangential caustic has a larger area than the radial caustic at all but the lowest eccentricities; there is a narrow eccentricity interval in which the probability of five-image configurations just exceeds that of three-image configurations. For higher values of $\kappa_{\text{s}}$: with increasing $e$ the tangential critical curve shrinks vertically and horizontally and its area drops nearly linearly with increasing ellipticity; the radial critical curve is elongated horizontally and its dimensions decrease monotonically with $e$; the radial caustic has a larger area than the tangential caustic at all but the highest eccentricities; the probability of five-image configurations never exceeds that of three-image configurations. Radial and tangential critical curves and caustics exist for any lens-parameter combination $(\kappa_{\text{s}},e)$. In the limit $e\to 1$ the radial critical curve and caustic both shrink to the point at the origin, while the tangential critical curve and caustic both shrink to a line segment extending from the origin along the major axis out to points with convergence $\kappa=1/2$.
@@
In Sect. 4.5 we investigated the properties of individual images occurring in the image plane of the ellipsoidal NFW lens for different parameter combinations. First, we used the technique of convergence–shear diagrams (Karamazov & Heyrovský, 2022) in Fig.s 1516 to identify the magnification and flattening of images occurring along concentric ellipses parameterized by their semi-major axis $a$ in the image plane. Second, we mapped the orientation of the image distortions in Fig. 17. Along the major axis, outside the unit-convergence ellipse images are always distorted vertically (i.e., tangentially) and inside they are always distorted horizontally (i.e., radially). However, along the minor axis the situation is more complicated: images are generally distorted horizontally except in the segments between the ellipse and the zero-shear points. As seen in the second panel of Fig. 17, images along the minor axis may be distorted radially at the tangential critical curve and tangentially at the radial critical curve.
@@
The methods that were employed and developed in this work to illustrate the properties of the ellipsoidal NFW lens may be applied in a similar manner to elucidate the lensing characteristics of other models of gravitational lenses.
@@ Appendix A Transformations between reference frames
We may now substitute the proper-frame coordinates from Eq. (71) into Eq. (2) for the semi-major axis $\hat{a}$ of a constant-density ellipsoid passing through a given point. In the latter equation, $e_{1}$ is the eccentricity of elliptical density contours in planes perpendicular to the minor axis $\hat{x}_{3}$, $e_{2}$ is the eccentricity of elliptical density contours in planes perpendicular to the median axis $\hat{x}_{2}$, and $10\leq e_{1}\leq e_{2}<$ 1. The equation can be solved in terms of the observer’s coordinates to yield the line-of-sight intersections with the ellipsoid at a given plane-of-the-sky position,
@@
In Fig. 1 their orientation is indicated by the axes $x_{1}$ and $x_{2}$. These are plane-of-the-sky coordinates rotated to the principal axes of the two-dimensional projected ellipse, which can be seen by setting $x_{3+}^{\prime}=x_{3-}^{\prime}$ and using Eq. (72). The semi-major axis of the ellipse is $\hat{a}\sqrt{\lambda_{1}}$ and it is oriented along the $\tilde{x}_{1}$ axis; the semi-minor axis is $\hat{a}\sqrt{\lambda_{2}}$ and it is oriented along the $\tilde{x}_{2}$ axis. Note that the first matrix in Eq. (75) including the square-root factor reduces to the identity matrix for any ellipsoid with $\psi=k\,\pi/2$ or $\vartheta=\pi/2$, or for an oblate ellipsoid ($e_{1}=0$) in an arbitrary orientation. In these cases Eq. (75) purely corresponds to the Eulerian rotation by $\varphi$ about $x_{3}^{\prime}$.
@@
The formulae presented above may be particularly useful when combining multiple halos with different spatial orientations, such as halos of galaxies in a group, or subhalos of individual galaxies in a galaxy-cluster halo.
@@ Appendix B Lensing quantities at the foci of the $a=1$ ellipse
The expressions for the deflection angle derived in Sect. 4.2 and for the shear derived in Sect. 4.3 seem to be undefined at $\bm{x}=(\pm e,0)$, the foci of an ellipse with semi-major axis equal to the projected scale semi-major axis of the halo ($a=1$). Nevertheless, all lensing quantities are finite and continuous at these points, which can be shown by computing the respective limits.
@@
The shear components from Eq. (35) and Eq. (36) attain the values
@@
The expressions presented above can be used in case of running into numerical problems at the foci.
@@ Appendix C Image-plane analysis of critical curves and caustics
We present here a new technique for analyzing and visualizing the properties of critical curves and caustics of gravitational lens models. The method is related to the image-plane analysis of lenses consisting purely of a set of point masses introduced by Daněk & Heyrovský (2015a), and applied to the analysis of triple lenses by Daněk & Heyrovský (2015b, 2019). In this appendix we thus use the term “image plane” instead of the synonymous “lens plane” used in the rest of the paper. We illustrate the technique on the example of the ellipsoidal NFW lens and demonstrate its usefulness for identifying the parameter-space boundaries in Fig. 8 and Fig. 10.
@@ C.1 Critical curves as Jacobian-factor contours
because $\kappa_{\text{s}}$ appears purely as a multiplicative factor in the convergence and in the shear, as seen from Eq. (21) and Eq. (38), respectively.
@@ C.2 Inflection points of critical curves
The concept of visualizing an entire one-parameter class of critical curves in a single tangential or radial contour plot can be developed further by tracking specific properties of the critical curves to identify parameter-space boundaries such as those plotted in Fig. 8 and Fig. 10. Here we demonstrate this approach on the occurrence of inflection points along the critical curves.
@@
The derivatives of the convergence and of the shear components can be computed from Eq. (21) and Eqs. (35)–(36); the second derivatives can be computed analogously. All derivatives of $j_{\text{T}}(\bm{x})$ are directly proportional to the convergence parameter $\kappa_{\text{s}}$. The left-hand side of Eq. (95) is thus clearly a homogeneous function of $\kappa_{\text{s}}$. Division by $\kappa_{\text{s}}^{3}$ shows that the inflection curve is independent of $\kappa_{\text{s}}$. We can therefore plot it over the contours of $j_{\text{T}}$ for a given eccentricity $e$; its intersections with the contours then indicate the locations of inflection points along tangential critical curves for all values of $\kappa_{\text{s}}$.
@@ C.3 Cusp number
Note that this zero-tangent-vector condition is automatically satisfied also by umbilic points, for which the entire Jacobi matrix is zero. However, the umbilic points on the caustic are not cusps, so we have to sort them out from the solutions of Eq. (101).
@@
Note that the two equations are equivalent off the major and minor axis, which can be seen for example by multiplying the first equation by $(\gamma+\gamma_{1})/\gamma_{2}$. Along the axes, $\gamma_{2}=0$ and one of the equations reduces to the trivial $0=0$ while the other requires one partial derivative of $j_{\text{T}}$ to be zero. Hence, for finding all cusps both equations should be preserved.
@@
However, the set of Eq. (102) and Eq. (103) is better suited for numerical solution, since solving Eq. (104) amounts to the more challenging task of finding a curve minimizing the non-negative function on the left-hand side. Finally, note that expressing the shear components in terms of shear and phase using Eq. (33) and Eq. (34) leads to a more geometric formulation of the tangential-caustic cusp condition,
@@
We call the curve defined by the common solution of Eq. (102) and Eq. (103) the tangential cusp curve. Division of both equations by $\kappa_{\text{s}}^{2}$ shows that the tangential cusp curve is independent of $\kappa_{\text{s}}$. In a similar manner as we did in the case of the inflection curve in Sect. C.2, we can thus plot it over the contours of $j_{\text{T}}$ for a given eccentricity $e$ and its intersections with the contours will indicate the locations of cusp images along tangential critical curves for all values of $\kappa_{\text{s}}$.
@@
Just as in the tangential case, the set of Eq. (106) and Eq. (107) is better suited for numerical solution. Expressing the shear components in terms of shear and phase using Eq. (33) and Eq. (34) leads to a more geometric formulation of the radial-caustic cusp condition
@@
We call the curve defined by the common solution of Eq. (106) and Eq. (107) the radial cusp curve. This curve is also independent of $\kappa_{\text{s}}$, and we can thus plot it over the contours of $j_{\text{R}}$ for a given eccentricity $e$, so that its intersections with the contours will indicate the locations of cusp images along radial critical curves for all values of $\kappa_{\text{s}}$.

### v2

@@ 5.1 Relevance of high-eccentricity results
The high-eccentricity and high-convergence-parameter region of parameter space mapped in Fig. 20, described in Sect. 4.4.3 and Appendix D, is the domain not only of the seven-image lens configurations, but also of bulge-shaped critical curves and other peculiarities not occurring at lower eccentricities. These properties of the ellipsoidal NFW lens could not be discovered from low-eccentricity approximations of the lens model, such as those derived from an elliptically distorted lens potential (e.g., Gomer et al., 2023). On the other hand, the astrophysical relevance of such parameter values should be considered.
@@ 5.2 Role of source redshift
The convergence parameter $\kappa_{\text{s}}$ is directly proportional to the angular-diameter-distance ratio $D_{\text{ls}}/D_{\text{s}}$, which increases with source redshift at a rate that depends on the cosmological model (see, e.g., Umetsu, 2020). While some of the presented results such as those in Figs. 26 or 12 are independent of $\kappa_{\text{s}}$ and thus of the source redshift, other are presented for specific values of $\kappa_{\text{s}}$, such as those in Figs. 7, 9, 14, or 1617. In the parameter-space maps in Figs. 8 and 20, different source-plane distances behind a given lens correspond to points along horizontal line segments. These may intersect the plotted boundaries, leading to significant changes in the critical curves and caustics as a function of source redshift.
The dependence of the convergence–shear diagrams on the source redshift is illustrated by the dependence on $\kappa_{\text{s}}$ in Fig. 15. The Jacobian-factor contour plots described in Appendix C (with examples in Figs. 18 and 19) present an excellent visualization of the variation of the critical curves with source redshift. Since different values of $\kappa_{\text{s}}$ correspond to different nested contours in the same plot, the critical curves for a given range of source redshifts can be directly seen as the corresponding sequence of neighboring contours. For a given lens at a redshift $z_{\text{l}}$, the contours could just as well be labeled by the source redshift $z_{\text{s}}$, with $z_{\text{s}}=z_{\text{l}}$ corresponding to the point at the origin of the plots and higher $z_{\text{s}}$ corresponding to larger contours.
@@ 5.3 Comparison with non-singular isothermal ellipsoid
The comparison in the previous paragraph is relevant to the low-convergence-parameter regime of the ellipsoidal NFW model discussed in Sect. 4.4.2 and mapped in the left panel of Fig. 8. The additional boundaries and lensing regimes appearing at higher $\kappa_{\text{s}}$, discussed in Sect. 4.4.3 and Appendix D and mapped in the right panel of Fig. 8 and in Fig. 20, do not occur in the NIE model. A more detailed comparison of these or other ellipsoidal models is beyond the scope of the present work.
5.4 Possible modifications of the model
@@
While the derived analytic formulae for the light deflection by the ellipsoidal NFW model present a large step toward a more astrophysically relevant model from the spherical or ellipsoidal-potential NFW models, structure-formation simulations indicate the potential further steps toward more realistic models. One would be the possible introduction of a core to eliminate the central density divergence (e.g., Salucci, 2019). Another would be to relax the homoeoidal symmetry of the model: simulations of dark matter halos such as those by Jing & Suto (2002) indicate that their nested constant-density surfaces do not preserve the same ellipsoidal shape. Rather, the axis ratios (here parameterized by the eccentricities $e_{1}$ and $e_{2}$) and axis orientations tend to vary from the center outward. However, such refinements would most probably stay in the realm of numerical modeling and their verification would likely require uniquely precise observational data.
@@ 6 Summary
We applied the described approach in Sect. 4 to study gravitational lensing by a triaxially ellipsoidal dark-matter halo with a NFW density profile (Navarro et al., 1996). The form of the convergence profile given by Eq. (21) as a function of semi-major axis is analogous to the convergence of a spherical NFW halo. However, the convergence parameter depends on the halo shape and orientation, as seen in Eq. (22). The main results of the present work are the analytic formulae derived in Sect. 4.2. The deflection angle for the studied ellipsoidal NFW halo is given by Eq. (25) in terms of the complex scattering function $I$. The components of the reduced deflection angle $\bm{\alpha}_{0}$ are presented in Eqs. (30)–(31) in terms of real functions. These exact closed-form expressions are valid for arbitrary eccentricity $e\in[0,1]$ of the surface-density profile.
@@
The variation of the linear dimensions and enclosed areas of the critical curves and caustics with eccentricity $e$ is demonstrated in Figs. 11 and 13, respectively. A comparison of the top and bottom rows shows there is a substantial change from a low-convergence to a higher-convergence regime. For low values of $\kappa_{\text{s}}$: with increasing $e$ the tangential critical curve grows prominently horizontally and its area initially declines before increasing to a high-eccentricity peak; the radial critical curve is elongated vertically and its vertical extent initially grows with $e$ before reaching the umbilic transition; the tangential caustic has a larger area than the radial caustic at all but the lowest eccentricities; there is a narrow eccentricity interval in which the probability of five-image configurations just exceeds that of three-image configurations. For higher values of $\kappa_{\text{s}}$: with increasing $e$ the tangential critical curve shrinks vertically and horizontally and its area drops nearly linearly with increasing ellipticity; the radial critical curve is elongated horizontally and its dimensions decrease monotonically with $e$; the radial caustic has a larger area than the tangential caustic at all but the highest eccentricities; the probability of five-image configurations never exceeds that of three-image configurations. Radial and tangential critical curves and caustics exist for any lens-parameter combination $(\kappa_{\text{s}},e)$. In the limit $e\to 1$ the radial critical curve and caustic both shrink to the point at the origin, while the tangential critical curve and caustic both shrink to a line segment extending from the origin along the major axis out to points with convergence $\kappa=1/2$.
@@
In Sect. 4.5 we investigated the properties of individual images occurring in the image plane of the ellipsoidal NFW lens for different parameter combinations. First, we used the technique of convergence–shear diagrams (Karamazov & Heyrovský, 2022) in Figs. 14 and 15 to identify the magnification and flattening of images occurring along concentric ellipses parameterized by their semi-major axis $a$ in the image plane. Second, we mapped the orientation of the image distortions in Fig. 16. Along the major axis, outside the unit-convergence ellipse images are always distorted vertically (i.e., tangentially) and inside they are always distorted horizontally (i.e., radially). However, along the minor axis the situation is more complicated: images are generally distorted horizontally except in the segments between the ellipse and the zero-shear points. As seen in the second panel of Fig. 16, images along the minor axis may be distorted radially at the tangential critical curve and tangentially at the radial critical curve.
@@
In this work we developed and employed a range of methods to illustrate the properties of the ellipsoidal NFW lens. These techniques may be applied in a similar manner to elucidate the lensing characteristics of other models of gravitational lenses.
@@ Appendix A Transformations between reference frames
We may now substitute the principal-frame coordinates from Eq. (71) into Eq. (2) for the semi-major axis $\hat{a}$ of a constant-density ellipsoid passing through a given point. In the latter equation, $e_{1}$ is the eccentricity of elliptical density contours in planes perpendicular to the minor axis $\hat{x}_{3}$, $e_{2}$ is the eccentricity of elliptical density contours in planes perpendicular to the median axis $\hat{x}_{2}$, and $10\leq e_{1}\leq e_{2}<$ 1. The equation can be solved in terms of the observer’s coordinates to yield the line-of-sight intersections with the ellipsoid at a given plane-of-the-sky position,
@@
In Fig. 1 their orientation is indicated by the axes $x_{1}$ and $x_{2}$. These are plane-of-the-sky coordinates rotated to the principal axes of the two-dimensional projected ellipse, which can be seen by setting $x_{3+}^{\prime}=x_{3-}^{\prime}$ and using Eq. (72). The semi-major axis of the ellipse is $\hat{a}\sqrt{\lambda_{1}}$ and it is oriented along the $\tilde{x}_{1}$ axis; the semi-minor axis is $\hat{a}\sqrt{\lambda_{2}}$ and it is oriented along the $\tilde{x}_{2}$ axis. We note that the first matrix in Eq. (75) including the square-root factor reduces to the identity matrix for any ellipsoid with $\psi=k\,\pi/2$ or $\vartheta=\pi/2$, or for an oblate ellipsoid ($e_{1}=0$) in an arbitrary orientation. In these cases Eq. (75) purely corresponds to the Eulerian rotation by $\varphi$ about $x_{3}^{\prime}$.
@@
The formulae presented above may be particularly useful when combining multiple halos with different spatial orientations. Examples of such systems include the halos of galaxies in a group, or the subhalos of individual galaxies in a galaxy-cluster halo.
@@ Appendix B Lensing quantities at the foci of the $a=1$ ellipse
The expressions for the deflection angle derived in Sect. 4.2 and for the shear derived in Sect. 4.3 seem to be undefined at $\bm{x}=(\pm e,0)$, the foci of an ellipse with semi-major axis equal to the projected scale semi-major axis of the halo ($a=1$). Nevertheless, all lensing quantities are finite and continuous at these points, which can be shown by computing the respective limits. We present here the obtained expressions which can be used in case of running into numerical problems at the foci.
@@
The shear components from Eqs. (35) and (36) attain the values
@@ Appendix C Image-plane analysis of critical curves and caustics
We present here a new technique for analyzing and visualizing the properties of critical curves and caustics of gravitational lens models. The method is related to the image-plane analysis of lenses consisting purely of a set of point masses introduced by Daněk & Heyrovský (2015a), and applied to the analysis of triple lenses by Daněk & Heyrovský (2015b, 2019). In this appendix we thus use the term “image plane” instead of the synonymous “lens plane” used in the rest of the paper. We illustrate the technique on the example of the ellipsoidal NFW lens and demonstrate its usefulness for identifying the parameter-space boundaries in Figs. 8 and 20.
@@ C.1 Critical curves as Jacobian-factor contours
because $\kappa_{\text{s}}$ appears purely as a multiplicative factor in the convergence and in the shear, as seen from Eqs. (21) and (38), respectively.
@@ C.2 Inflection points of critical curves
The concept of visualizing an entire one-parameter class of critical curves in a single tangential or radial contour plot can be developed further by tracking specific properties of the critical curves to identify parameter-space boundaries such as those plotted in Figs. 8 and 20. Here we demonstrate this approach on the occurrence of inflection points along the critical curves.
@@
The derivatives of the convergence and of the shear components can be computed from Eqs. (21), (35), and (36); the second derivatives can be computed analogously. All derivatives of $j_{\text{T}}(\bm{x})$ are directly proportional to the convergence parameter $\kappa_{\text{s}}$. The left-hand side of Eq. (95) is thus clearly a homogeneous function of $\kappa_{\text{s}}$. Division by $\kappa_{\text{s}}^{3}$ shows that the inflection curve is independent of $\kappa_{\text{s}}$. We can therefore plot it over the contours of $j_{\text{T}}$ for a given eccentricity $e$; its intersections with the contours then indicate the locations of inflection points along tangential critical curves for all values of $\kappa_{\text{s}}$.
@@ C.3 Cusp number
We note that this zero-tangent-vector condition is automatically satisfied also by umbilic points, for which the entire Jacobi matrix is zero. However, the umbilic points on the caustic are not cusps, so we have to sort them out from the solutions of Eq. (101).
@@
We note that the two equations are equivalent off the major and minor axis, which can be seen for example by multiplying the first equation by $(\gamma+\gamma_{1})/\gamma_{2}$. Along the axes, $\gamma_{2}=0$ and one of the equations reduces to the trivial $0=0$ while the other requires one partial derivative of $j_{\text{T}}$ to be zero. Hence, for finding all cusps both equations should be preserved.
@@
However, the set of Eqs. (102) and (103) is better suited for numerical solution, since solving Eq. (104) amounts to the more challenging task of finding a curve minimizing the non-negative function on the left-hand side. Finally, we note that expressing the shear components in terms of shear and phase using Eqs. (33) and (34) leads to a more geometric formulation of the tangential-caustic cusp condition,
@@
We call the curve defined by the common solution of Eqs. (102) and (103) the tangential cusp curve. Division of both equations by $\kappa_{\text{s}}^{2}$ shows that the tangential cusp curve is independent of $\kappa_{\text{s}}$. In a similar manner as we did in the case of the inflection curve in Sect. C.2, we can thus plot it over the contours of $j_{\text{T}}$ for a given eccentricity $e$ and its intersections with the contours will indicate the locations of cusp images along tangential critical curves for all values of $\kappa_{\text{s}}$.
@@
Just as in the tangential case, the set of Eqs. (106) and (107) is better suited for numerical solution. Expressing the shear components in terms of shear and phase using Eqs. (33) and (34) leads to a more geometric formulation of the radial-caustic cusp condition
@@
We call the curve defined by the common solution of Eqs. (106) and (107) the radial cusp curve. This curve is also independent of $\kappa_{\text{s}}$, and we can thus plot it over the contours of $j_{\text{R}}$ for a given eccentricity $e$, so that its intersections with the contours will indicate the locations of cusp images along radial critical curves for all values of $\kappa_{\text{s}}$.